OpenStudy (elleblythe):

How to convert this to its polar form: z^4+8√3=8i

OpenStudy (dan815):

hi

OpenStudy (dan815):

z^4=-8r3+8i

OpenStudy (dan815):

z^4=re^itheta r=mag(-8r3+8i) theta = taninv(8/-8root3)

OpenStudy (dan815):

then u can take qad root

OpenStudy (dan815):

by dividing theta and qadroot of mag

OpenStudy (dan815):